line parallel to the line y = 2x + 5 and through the point (2,-3) whats the answer
Answer by mananth(16949)

Find the slope of this line
y = 2 x + 5
Divide by 1
y = 2 x + 5
Compare this equation with y=mx+b
slope m = 2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 2
m= 2 ,point (2,-3)
Find b by plugging the values of m & the point in
y=mx+b
-3 = 4.00 + b
b= -7
m= 2
Plug value of the slope and b in y = mx +b
The required equation is y = 2 x -7
y = 2 x -7
m.ananth@hotmail.ca